the mole fraction of H2so4(sulfuric acid)in a soln(solution)containg 98% h2so4 by mass

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
Density = 2.08 g/cc (assuming)
Weight of 1 litre of solution = 2.08 *1000 = 2080g
gram molecular weight of H2SO4 = 98g
Moles of H2SO4 = (2080-1000)/98 = 11.02 moles( 1 litre = 1000g and density of water=1)
moles of water = 1000/18 = 55.56
Mole fraction = 11.02/11.02+55.56 = 0.166.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)